Please Help me answer this homework promblem on equations. a one year old killer whale is 290 cm long the whale grows 3cm per month how many years will it take until the whale is 470 cm long? Please help thank you
Answer by jim_thompson5910(35256)

Let x = months, y = length

Equation is: y = 3x + 290

You get this by adding the initial length of 290 cm to the variable length 3x cm. For example, in month zero (the start), the length is y = 3(0)+290 = 290 and in one month, the length is y = 3(1) + 290 = 293


From here, plug in y = 470 and solve for x:


y = 3x + 290

470 = 3x + 290

470-290 = 3x

180 = 3x

180/3 = x

60 = x

x = 60

So in 60 months, or 60/12 = 5 years, the whale will be 470 cm long